# UrbanVLP: A Multi-Granularity Vision-Language Pre-Trained Foundation Model for Urban Indicator Prediction

### v1

@@ 4.1.2. Downstream Task Dataset
In our research, the Downstream Task Dataset plays a crucial role in gauging the practical implications of our multimodal framework. We collect six representative indicators: Carbon Emission, Population, GDP, Night Light, House Price, and POI. Each indicator illuminates a significant dimension of urban vitality and sustainability. The detailed introduction of each downstream indicator is provided in the Appendix A.
Table 2. Urban indicators prediction results in four datasets. The best results are in bold and the second-best results are underlined.
Methods UrbanVLP StructualUrban (Li et al., 2022c) Vision-LSTM (Huang et al., 2023a) UrbanCLIP-SV (Yan et al., 2023) UrbanCLIP (Yan et al., 2023) PG-SimCLR (Xi et al., 2022) ResNet-18 (He et al., 2016) Autoencoder (Kramer, 1991)
Metric $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE
Beijing Carbon 0.787 0.457 0.353 0.745 0.500 0.397 0.692 0.502 0.421 0.504 0.697 0.536 0.716 0.528 0.528 0.430 0.797 0.632 0.383 0.821 0.598 0.294 0.851 0.572
Population 0.725 0.513 0.404 0.699 0.536 0.422 0.649 0.583 0.511 0.447 0.727 0.574 0.661 0.569 0.452 0.476 1.228 0.963 0.251 0.874 0.657 0.164 0.918 0.673
GDP 0.586 0.650 0.416 0.560 0.670 0.418 0.536 0.690 0.454 0.208 0.899 0.553 0.555 0.674 0.439 0.270 1.679 1.067 0.242 1.253 0.733 0.165 1.521 0.838
Night Light 0.531 0.629 0.394 0.471 0.668 0.447 0.409 0.712 0.466 0.352 0.733 0.476 0.420 0.700 0.457 0.367 0.728 0.404 0.337 0.740 0.581 0.276 0.784 0.643
House Price 0.506 0.640 0.479 0.494 0.647 0.481 0.470 0.655 0.493 0.453 0.672 0.512 0.503 0.641 0.482 0.341 0.718 0.555 0.284 0.721 0.576 0.332 0.714 0.576
POI 0.319 0.706 0.383 0.277 0.712 0.385 0.236 0.718 0.388 0.227 0.724 0.410 0.253 0.731 0.388 - - - 0.209 0.747 0.415 0.164 0.759 0.427
Shanghai Carbon 0.716 0.529 0.392 0.677 0.556 0.431 0.568 0.613 0.458 0.574 0.473 0.453 0.671 0.569 0.426 0.270 0.742 0.551 0.243 0.763 0.528 0.230 0.771 0.532
Population 0.593 0.607 0.471 0.548 0.630 0.486 0.521 0.657 0.518 0.402 0.793 0.590 0.536 0.648 0.507 0.288 1.051 0.831 0.274 0.833 0.632 0.166 0.898 0.742
GDP 0.310 0.816 0.595 0.305 0.841 0.613 0.304 1.012 0.664 0.315 1.214 0.741 0.326 0.807 0.587 0.270 1.679 1.067 0.241 1.237 0.740 0.197 1.511 0.844
Night Light 0.457 0.667 0.459 0.412 0.689 0.494 0.355 0.731 0.546 0.276 0.763 0.554 0.387 0.709 0.511 0.228 0.596 0.466 0.202 0.789 0.573 0.164 0.810 0.617
POI 0.413 0.673 0.363 0.387 0.681 0.375 0.247 0.699 0.389 0.243 0.707 0.399 0.269 0.704 0.394 - - - 0.221 0.723 0.411 0.186 0.729 0.425
Guangzhou Carbon 0.698 0.514 0.385 0.654 0.550 0.506 0.551 0.564 0.462 0.423 0.520 0.478 0.585 0.603 0.444 0.413 0.951 0.716 0.362 0.673 0.536 0.254 0.733 0.570
Population 0.665 0.565 0.441 0.638 0.586 0.453 0.617 0.615 0.506 0.497 0.685 0.612 0.633 0.592 0.467 0.294 1.046 2.112 0.249 0.834 0.678 0.195 0.928 0.753
GDP 0.436 0.764 0.541 0.419 0.791 0.549 0.405 1.011 0.625 0.321 1.004 0.689 0.440 0.762 0.546 0.263 2.162 1.553 0.234 1.147 0.732 0.176 1.463 0.811
Night Light 0.577 0.573 0.418 0.534 0.596 0.442 0.405 1.011 0.625 0.335 0.672 0.514 0.483 0.633 0.478 0.436 0.549 0.427 0.225 0.720 0.562 0.176 0.798 0.620
POI 0.396 0.752 0.344 0.325 0.771 0.364 0.143 0.822 0.411 0.136 0.823 0.415 0.150 0.819 0.408 - - - 0.108 0.920 0.433 0.084 0.956 0.469
Shenzhen Carbon 0.659 0.568 0.418 0.613 0.570 0.434 0.564 0.576 0.471 0.483 0.586 0.488 0.562 0.571 0.483 0.234 0.975 0.746 0.207 0.735 0.597 0.189 0.887 0.634
Population 0.790 0.448 0.343 0.724 0.476 0.384 0.624 0.562 0.471 0.589 0.604 0.483 0.727 0.510 0.392 0.294 1.046 2.112 0.264 0.871 0.694 0.175 0.921 0.772
GDP 0.532 0.676 0.448 0.489 0.695 0.471 0.462 0.720 0.517 0.433 0.754 0.577 0.508 0.693 0.464 0.294 1.046 2.112 0.182 1.245 0.836 0.119 1.733 0.884
Night Light 0.457 0.667 0.459 0.421 0.694 0.502 0.404 0.702 0.508 0.320 0.717 0.531 0.387 0.709 0.511 0.454 0.199 0.158 0.243 0.726 0.543 0.166 0.758 0.608
POI 0.461 0.759 0.347 0.321 0.780 0.374 0.254 0.795 0.396 0.147 0.849 0.440 0.185 0.838 0.427 - - - 0.126 0.968 0.469 0.092 0.973 0.479
$1^{\text{st}}$ount 53 0 0 1 5 3 0 0
Table 3. Ablation study on UrbanSV Beijing dataset. The best results are in bold.
Model Carbon Population GDP Night Light House Price POI
$\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE
w/o SV branch 0.754 0.491 0.387 0.690 0.545 0.428 0.547 0.680 0.433 0.468 0.670 0.445 0.520 0.635 0.465 0.263 0.714 0.387
w/o ST branch 0.518 0.687 0.512 0.465 0.715 0.559 0.229 0.887 0.543 0.423 0.698 0.423 0.295 0.764 0.575 0.265 0.716 0.390
w/o LE branch 0.776 0.469 0.358 0.717 0.520 0.414 0.589 0.652 0.409 0.518 0.633 0.400 0.508 0.639 0.472 0.301 0.710 0.385
w/o SV&ST branch 0.716 0.528 0.324 0.626 0.598 0.425 0.258 0.871 0.511 0.482 0.661 0.402 0.493 0.654 0.513 0.259 0.721 0.395
UrbanVLP 0.788 0.455 0.455 0.720 0.517 0.410 0.585 0.651 0.420 0.523 0.634 0.398 0.523 0.634 0.407 0.319 0.706 0.383
4.2. Experimental Setup
4.2.1. Baselines
We compare our method with 7 baselines in the field of urban imagery-based socioeconomic prediction. The single granularity imagery-based urban region representation learning methods include: AutoEncoder (Kramer, 1991), ResNet-18 (He et al., 2016), PG-SimCLR (Xi et al., 2022), UrbanCLIP (Yan et al., 2023), UrbanCLIP-SV (Yan et al., 2023). We also apply multiple baselines for multi-granularity urban imagery-based representation learning: Vision-LSTM (Huang et al., 2023b) and StructualUrban
(Li et al., 2022c)
. The in-depth overview of each baseline method is presented in the Appendix B.
4.2.2. Metrics and Implementation
@@
The performance of our predictions is evaluated using a trio of standard metrics: the coefficient of determination denoted as $R^{2}$, the root mean squared error (RMSE), and the mean absolute error (MAE). An increase in $R^{2}$, along with a reduction in RMSE and MAE values, is indicative of enhanced model accuracy. Adam optimizer is chosen to minimize the training loss during parameter learning. A grid search on hyperparameters is conducted, where search ranges for learning rate and batch size are set as 2e-6, 2e-5, 2e-4, 2e-3, 2e-2 and 4, 8, 16, 32, 64, respectively. The data augmentation strategy follows the setting of (Radford et al., 2021). The weight hyperparameter $\alpha$ and $\beta$ of $\mathcal{L}_{Total}$ are set to 0.5, 0.5 respectively. We run all the models on NVIDIA A800 GPUs.
@@
We implement the aggregation method $Aggr$ through MLP layers. To ensure a fixed length for network processing, we set the max street-view sequence length within one as 25, and sequences shorter than that are achieved by padding with zeros. The fusion method is Addition. We utilize the current state-of-the-art GeoCLIP (Vivanco et al., 2023) as our location encoder. For the semantic segmentation model to decouple visual elements with street-view images, the architecture for the semantic model is ResNet18dilated + PPM_deepsup. The original dataset contains 150 categories, from which we select 38 categories and group them into 13 street-view features following. (Fan et al., 2023b). The detailed categories can be seen in Appendix C.
4.3. RQ1: Performance Evaluation
To further validate the effectiveness of the proposed UrbanVLP, we conduct comparisons with existing state-of-the-art methods using all datasets. The comprehensive results are presented in Table 2. As we can see, UrbanVLP significantly surpasses the performance of baseline methods employing both single-granularity and multi-granularity imagery, achieving a substantial margin of improvement. It outperforms the best baseline, StructualUrban (Li et al., 2022c), by 4.2%, 2.6%, 2.6%, 6.0%, 1.2% and 4.7% in terms of $R^{2}$ for all 6 indicators: Carbon, Population, GDP, Night Light, House Price and POI in Beijing. The similar performance improvement can also be seen in the dataset of the other three cities. Besides, the average performance gain of UrbanVLP on RMSE and MAE in Beijing data are 2.3% and 2.0%, respectively. In order to provide a more intuitive display, we have created radar Image in Figure 2 to visualize the performance of two cities, Beijing and Shenzhen. The results further prove the effectiveness of our multi-granularity framework for urban region representation learning.
It is evident that pipelines leveraging multi-granularity approaches typically demonstrate superior performance compared to those employing a single granularity level. This enhancement can be rationalized by the incorporation of additional fine-grained information derived from street-view imagery. UrbanVLP exhibits superior performance over StructuralUrban (Li et al., 2022c) and Vision-LSTM (Huang et al., 2023b), illustrating the efficacy of textual mobility, which enhances interpretability. Concurrently, it is observed that UrbanCLIP-SV significantly underperforms compared to UrbanCLIP, demonstrating that street-view imagery, while rich in fine-grained details, lacks macroscopic perspective, leading to suboptimal outcomes when used in isolation. This also underscores the irreplaceability of satellite imagery as a macroscopic visual modality.
4.4. RQ2: Ablation Studies
We conduct ablation studies to investigate the effectiveness of different components in UrbanVLP on UrbanSV Beijing dataset, including the Satellite branch (ST branch), Street-View branch (SV branch), and Location Encoding branch (LE branch). The results are demonstrated in Table 3. It can be observed that the incorporation of street-view imagery results in an average improvement of 3.75% in terms of $R^{2}$. Furthermore, while street-view imagery contributes fine-grained information, its individual impact is still outstripped by that of satellite imagery, aligning with the performance differential between UrbanCLIP and UrbanCLIP-SV in Section 4.3. The role of location coordinates has been underexplored in previous work. Although not as impactful as visual modalities, their inclusion has yielded an average improvement of 0.5% in terms of $R^{2}$.
@@ 4.4. RQ2: Ablation Studies
In our work, we leverage ShareGPT4V (Chen et al., 2023a) for text description generation. Here we conduct experiments on current open-sourced LMM models, and compute the PerceptionScore of them. The results are presented in Table LABEL:table:lmms. As we can see, ShareGPT4V achieves the best performance in terms of PerceptionScore, followed by LLaMA-Adapter V2, which are used as the Image-to-Text models (Yan et al., 2023).
Table 4. Perception score of representative LMMs.
@@
LLaVA-v1.5 (Liu et al., 2023a) 0.674
mPLUG-Owl (Ye et al., 2023) 0.662
LLaMA-Adapter V2 (Gao et al., 2023) 0.698
MiniGPTv2 (Chen et al., 2023b) 0.693
ShareGPT4V (Chen et al., 2023a) 0.714
Figure 5. Illustration of the quality of our generated description example. The prompt for (c) is :”From scale 1 to 10, decide how similar the image is to the description”, whereas the prompt for (d) is: ”Generate a panoramic street-view image based on the description”.
Figure 6. Perception Score distribution on UrbanSV dataset for four representative cities.
4.5. RQ3: Qualitative Analysis
To visually illustrate the quality of the generated description, we present an example in Figure 5. The generated description of the street-view image (a) is depicted in (b). Subsequently, we employ GPT4V (Yang et al., 2023) to assess the quality score of the generated description, which is rated as 7 out of 10, indicating the effectiveness of the generated text. Furthermore, GPT4V also provides specific areas for improvement in (c). In (d), we utilize GPT4V to generate an image based on the description in (b). It is evident that the generated image bears a resemblance to the original one. In order to ascertain the efficacy of ShareGPT4V, we computed the distribution of PerceptionScores for descriptions generated by ShareGPT4V, as depicted in Figure 6. Notably, only less than 0.02% of image-text pairs scored below 0.5, thereby evidencing the high quality of the descriptions produced.
Figure 7. Examples of predictions from UrbanCLIP and our proposed UrbanVLP.
Here we also show some predicted results in Figure 7 on UrbanSV Beijing dataset. As we can see, satellite image (a) and (b), despite sharing similar layouts, (a) encompasses residential and campus areas, while (b) pertains to industrial zones. Their socio-economic characteristics are entirely different, which could be reflected by socio-enconomic indicators. The carbon emissions and GDP of campus and residential area are significantly lower than those of industrial park, whereas the disparities in population and housing prices are not as pronounced, possibly due to the unique characteristics of school district housing near educational institutions. Night light intensity is similar across the two area due to the fact that the majority of nocturnal economic activities within factories occur indoors, unlike in Central Business Districts (CBDs) where extensive illumination is generated. The same principle applies to school and residential area as well. As observed, the predictions made by UrbanCLIP do not effectively differentiate between the two, with the downstream metrics in Figure 7 (b) still inclined towards predicting a homogenized socio-economic functional area similar to that of (a). In contrast, the results from UrbanVLP are capable of distinctly distinguishing the socio-economic attributes of the two in terms of carbon emission, GDP, and house price.
5. Conclusion and future work
Urban region representation learning plays a significant role in helping us understand the patterns, structures, and dynamics of cities. UrbanVLP, for the first time, explores the differences between street-view and satellite images from a semantic granularity perspective, as well as their roles in modeling urban region representation. A text generation and calibration mechanism is also proposed to ensure high-quality description generation. It has achieved state-of-the-art results on the constructed UrbanSV and downstream dataset. In the future, the further research direction can be two-fold: Add additional modalities such as POIs to enhance information richness and enable the design of better model architectures such as MAE (He et al., 2022) to efficiently utilize existing data.
Appendix A Detailed introduction of downstream dataset
We provide the detailed introduction and data source of our downstream benchmark.
  • Carbon Emissions: Extracted from the Open-source Data Inventory for Anthropogenic CO2 (ODIAC), this dataset offers granular carbon emission figures in 2022 that are in alignment with the spatial resolution of our satellite imagery, which is one square kilometer per image. The unit is ton.
  • Population: Sourced from WorldPop, this data provides population distribution data in 2020. The unit is #citizens.
  • GDP: This data reflects the economic development situation in China in terms of Gross Domestic Product (GDP) figures to add an economic layer to our urban region analysis.
  • Night Light: Nighttime light can, to some extent, characterize the intensity of human activities, which is of great significance for urban development studies. we take the nightlight data from (Zhong et al., 2022) in 2020.
  • House Price: The price of housing is influenced by a variety of socio-economic factors. By observing changes in housing prices, it is possible to discern, to a certain extent, the state of economic development. We collect data from Anjuke (https://www.anjuke.com) in 2023.3. The unit is yuan/$m^{2}$. We take the average of housing prices within an urban region.
  • POI: The distribution of different types of POIs (Points of Interest) within urban areas can serve as a social indicator, as they reflect the social and economic structure, population activities, and lifestyles of the city. We collect POIs (location and category) from the web. The data has 14 categories: Dining and Cuisine, Leisure and Entertainment, Sports and Fitness, Business and Residential, Healthcare, Financial Institutions, Tourist Attractions, Lifestyle Services, Shopping and Consumption, Automobile Related, Hotel Accommodation, Transport Facilities, Science, Education and Culture, Companies and Enterprises. We take aggregation of the number of POIs of each category within each region.
We construct this benchmark to comprehensively evaluate the learned urban region representation. We uniformly partition the downstream dataset into train:val:test=6:2:2. These indicators constitute the downstream task dataset, covering economic (GDP, Night Light, House Price), social (Population, POI), and environmental (Carbon Emission) indicators. The data preprocessing setting adheres to (Yan et al., 2023) with refinement.
Appendix B In-depth overview of baseline methods
  • AutoEncoder (Kramer, 1991): An architectural framework for neural networks that is designed to extract representations from unlabeled satellite imagery with reconstruction loss.
  • ResNet-18 (He et al., 2016): The well-known residual neural network pre-trained on ImageNet (Deng et al., 2009), which can directly extract visual features of satellite imagery with the knowledge acquired from natural imagery.
  • PG-SimCLR (Xi et al., 2022): A contrastive learning framework which introduces geographic information (i.e., POI) into urban region representation learning. We don’t use PG-SimCLR to predict POI for a fair comparison in our study.
  • UrbanCLIP (Yan et al., 2023): A model that utilizes an image-text contrastive learning-based approach for learning robust representations from urban imagery, intending to capture the complexity and diversity of urban areas.
  • UrbanCLIP-SV (Yan et al., 2023): Original UrbanCLIP is pre-trained on satellite image-text pairs. Here we also verify its performance on our proposed UrbanSV dataset, utilizing the same aggregation method to acquire regional features.
  • Vision-LSTM (Huang et al., 2023b): A vision long short-term memory neural network which could process variable-length street-view image within each region. The integration of satellite and street-view imagery enables a more effective capture of comprehensive urban representation, achieving reliable recognition results.
  • StructualUrban(Li et al., 2022c): The graph-based framework for urban region profiling leveraging street segments as containers to adaptively fuse the features of multi-level urban images. We term this method StructualUrban for simplicity.
Appendix C Segmentation Ratio and Prompt Examples
As illustrated in Figure 8, we demonstrate examples of the segmentation ratio of street-view images in our UrbanSV dataset. The segmentation ratio is decoupled by pre-trained segmentation models and is used in both prompt construction and perception score computation. We group the segmentation categories into 13 categories following (Fan et al., 2023b): Light Vehicle; Façade; Window & Opening; Road; Sidewalk; Street Furniture; Greenery - Tree; Greenery - Grass & Shrubs; Sky; Nature. Figure 9 illustrates some exmaples of our prompts used for description generation.
Figure 8. Segmentation Ratio.
Figure 9. Prompt Examples.
Appendix D RQ4: Practicality
Figure 10. UrbanVLP system.

### v2

@@ B.2 Downstream Dataset & Tasks
In our research, the Downstream Dataset plays a crucial role in gauging the practical implications of our multimodal framework. We collect six representative urban indicators: Carbon Emission, Population, GDP, Night Light, House Price, and POI. Each indicator illuminates a significant dimension of urban vitality and sustainability. In this section, we provide the detailed introduction and data source of our downstream benchmark.
  • Carbon Emissions: Extracted from the Open-source Data Inventory for Anthropogenic CO2 (ODIAC), this dataset offers granular carbon emission figures in 2022 that are in alignment with the spatial resolution of our satellite imagery, which is one square kilometer per image. The emissions are quantified in tons.
  • Population: The quantum of population serves as an indicator of a region’s sociodemographic characteristics. Sourced from WorldPop (https://hub.worldpop.org), we gather population distribution data in 2020, to provide an in-depth portrayal of demographic patterns. The unit is #citizens.
  • GDP: This dataset encapsulates China’s economic development by incorporating Gross Domestic Product (GDP) figures, thereby infusing an economic dimension into our analysis of urban regions. The unit is million Chinese Yuan.
  • Night Light: Nighttime light can, to some extent, characterize the intensity of human activities, which is of great significance for urban development studies. we take the nightlight data from [87] in 2020.
  • House Price: The price of housing is is subject to the influence of numerous socio-economic factors. By observing fluctuations in housing prices, it is possible to discern, to a certain extent, the state of economic development. We collect data from Anjuke (https://www.anjuke.com) in March 2023 with yuan/$m^{2}$ unit. Individual house price represents a locational attribute. To extrapolate a regional-level characteristic, we take the average of housing prices across an urban region.
  • POI: The spatial distribution of various types of POIs (Points of Interest) within urban areas is a potent social indicator, as it reflects the underlying social and economic architecture, as well as the population’s activities and urban lifestyles. Consequently, to glean insights into these urban dynamics, we collect data of POIs from web sources, encompassing both their locations and categories. The collected data has 14 categories: Dining and Cuisine, Leisure and Entertainment, Sports and Fitness, Business and Residential, Healthcare, Financial Institutions, Tourist Attractions, Lifestyle Services, Shopping and Consumption, Automobile Related, Hotel Accommodation, Transport Facilities, Science, Education and Culture, Companies and Enterprises. To capture the unique attributes of each area, we aggregate the number of POIs by category within each region.
We construct this benchmark with the aim of conducting a thorough evaluation of the urban region representations that have been derived through our learning framework. We uniformly partition the downstream dataset into train:val:test=7:1:2. These indicators constitute the downstream task dataset, covering economic (GDP, Night Light, House Price), social (Population, POI), and environmental (Carbon Emission) indicators. The data preprocessing setting adheres to [74] with refinement. For some indicators data with relatively small differences in values, we have taken the logarithm of these values to facilitate training.
B.3 Downstream Dataset Analysis
Figure 12: Spatial distribution of Popultion, GDP and NightLight in Beijing.
Here we visualize the distribution of some downstream socio-economic indicator data on Beijing in Figure 12. The distribution of different indicators exhibits similar trends, indicating that Beijing has a clear ’center-periphery’ structure, with the core area being densely populated and economically developed, while the peripheral areas are relatively sparse.
Beyond geographical spatial differences, the concentration trends and variability of the data are critical perspectives that merit attention. In Figure 14 we demonstrate these patterns by illustrating the normalized values of different socio-economic indicators in Beijing. Outliers with extreme maximum and minimum values will be filtered out during our processing. As we can see, the values of Carbon Emission and NightLight are relatively concentrated, showing minimal variability. The distribution of GDP values is more widespread, while Population distribution exhibits a bimodal characteristic, reflecting the inherent diversity.
Figure 13: Data distribution of Carbon Emission, Popultion, GDP and NightLight in Beijing.
Figure 14: PerceptionScore distribution on $\mathtt{CityScape}$ dataset for 4 representative cities.
Appendix C In-depth Overview of Baseline Methods
  • ViT [15]: ViT is a successful paradigm of applying transformers to computer vision, where images are segmented into patches for fixed-length modeling. When sufficient data is available for pre-training, ViT’s performance can surpass that of CNNs [22]. In our experiment, we leverage ViT-B as the baseline for comparison.
  • PG-SimCLR [73]: A contrastive learning framework which introduces geographic information (i.e., POI) into urban region representation learning. In our study, we do not employ PG-SimCLR to predict POI to ensure a fair comparison.
  • UrbanCLIP [74]: A model that utilizes an image-text contrastive learning-based approach for learning robust representations from urban imagery, intending to capture the complexity and diversity of urban areas.
  • UrbanCLIP-SV [74]: Original UrbanCLIP is pretrained on satellite image-text pairs. Here we also verify its performance on our proposed $\mathtt{CityScape}$ dataset, utilizing the same aggregation method to acquire regional features.
  • Vision-LSTM [27]: A vision long short-term memory neural network which could process variable-length street-view image within each region. The integration of satellite and street-view imagery, along with trajectories, facilitates a more effective capture of comprehensive urban representations, yielding reliable recognition results.
  • StructuralUrban[41]: The graph-based framework for urban region profiling leveraging street segments as containers to adaptively fuse the features of multi-level urban images. We term this method StructuralUrban for simplicity.
We present a visual representation of the performance in the main paper Figure 4 to effectively demonstrate the comprehensive capabilities of our model across various downstream tasks. There is also a trend that models tend to perform better on environmental (Carbon Emission) and social (Population, POI) indicators than on economic (GDP, Night Light, House Price) indicators. This is possibly because environmental and social indicators usually have more distinct geographic distribution characteristics. For instance, neighboring regions may have similar climates, vegetation cover, and population densities, which are important factors for environmental and social indicators and can be effectively captured through spatial embedding. On the other hand, economic indicators are more complex and influenced by various factors, including policies, economic activities, and market sentiments, making them harder to model. Additionally, due to the multi-class nature of POI, models may require more complex structures to understand the underlying relationships between labels, thus resulting in lower performance.
Appendix D Implementation Details
@@
In our experiment, Adam optimizer is chosen to minimize the training loss during parameter learning. A grid search on hyperparameters is conducted, where search ranges for learning rate and batch size are set as 2e-6, 2e-5, 2e-4, 2e-3, 2e-2 and 4, 8, 16, 32, 64, respectively. The data augmentation strategy follows the setting of [61]. The weight hyperparameter $\alpha$ and $\beta$ of $\mathcal{L}_{Total}$ are set to 0.5, 0.5 respectively. We run all the models on NVIDIA A800 GPUs with PyTorch.
@@
We implement the aggregation method $\operatorname{Aggr}(\cdot)$ through MLP layers. The feature fusion method f  directly utilizes additive aggregation. To ensure a fixed length for network processing, we set the maximum street-view sequence length within one region as 25, and sequences shorter than that are achieved by padding with zeros. For the semantic segmentation model to decouple visual elements with street-view images, following  [17], the architecture for the semantic model is ResNet18dilated + PPM_deepsup. Initially, the dataset comprises 150 segmentation categories. However, in line with our requirements, we narrow down the selection to 38 categories, which we further group into 13 street-view features following [17]: Person; Bike; Heavy Vehicle; Light Vehicle; Facade; Window & Opening; Road; Sidewalk; Street Furniture; Greenery - Tree; Greenery - Grass & Shrubs; Sky; Nature.
Appendix E More Details for Modality Representation
For utilizing ViT as the visual encoder to process urban visual data, in particular, we reshape the input image $I\in\mathbb{R}^{H\times W\times 3}$ into a sequence of flattened 2D patches ${I}_{P}\in\mathbb{R}^{N\times({P^{2}}\cdot{C})}$, where C represents the number of channels, $(P,P)$ specifies the resolution of each image patch, and $N=HW/P^{2}$ denotes the number of patches, corresponding to the input sequence length for ViT. Notebly, the ViT utilizes a consistent latent vector size (denoted as $d$) across all its layers. Thus, we can obtain the outputs as the patch embeddings through a trainable linear projection. We also prepend a learnable embedding to the patch embedding sequence (i.e., $\mathbf{z}_{0}^{0}={I}_{cls}$). Besides, we add learnable position embedding $\mathbf{E}_{pos}$ to retain positional information. The process is as follows:
$\mathbf{z}_{0}=[{I}_{cls};{I}_{p}^{1}\mathbf{W};{I}_{p}^{2}\mathbf{W};\cdots;{I}_{p}^{N}\mathbf{W}]+\mathbf{E}_{pos},$ (8)
where $\mathbf{W}\in\mathbb{R}^{(P^{2}\cdot C)\times d}$ and $\mathbf{E}_{pos}\in\mathbb{R}^{(N+1)\times d}$ are learnable parameters. $\mathbf{z}_{0}$ represents the result of the linear projection, which is subsequently fed into the ViT.
Consistent with the standard Transformer encoder, ViT includes alternating layers of multi-head self-attention $\operatorname{MSA(\cdot)}$ operation [68]. Furthermore, layer normalization (LN) is adopted before every block, whereas residual connections are after every block:
$\mathbf{z}_{l}=\mathrm{MSA}(\mathrm{LN}(\mathbf{z}_{l-1}))+\mathbf{z}_{l-1},\quad l=1\ldots L.$ (9)
The text encoder follows a similar $\operatorname{MSA}$ mechanism of the visual encoder. The input text sequence is bracketed with $[SOS]$ and $[EOS]$ tokens, and the activation of the highest layer of Transformer at $[EOS]$ token is considered the global representation $\mathbf{z}_{T}$ of text.
Appendix F Segmentation Ratio and Prompt Examples
As illustrated in Figure 15(a) and (b), we demonstrate examples of the segmentation ratio of street-view images in our $\mathtt{CityScape}$ dataset. The segmentation ratio is decoupled by pretrained segmentation models and is used in both prompt construction and PerceptionScore computation. We group the segmentation categories into 13 categories following [17]: Person; Bike; Heavy Vehicle; Light Vehicle; Facade; Window & Opening; Road; Sidewalk; Street Furniture; Greenery - Tree; Greenery - Grass & Shrubs; Sky; Nature.
The prompt template used to generate street-view descriptions is formulated as follows: Analyze the street-view panoramic image in [City] in a comprehensive and detailed manner. The coordinate of the street-view image is [Longitude, Latitude]. The segmentation ratio of the street-view image is [Segment Information]. Figure 15(c) provides illustrations of some examples of our prompts used for description generation.
Figure 15: Prompt Examples.
Appendix G The Choice of Image-to-Text LMM
@@
In our work, we leverage ShareGPT4V [11] for text description generation. Here we conduct experiments on current open-sourced LMM models, and compute the PerceptionScore of them. The results are presented in Table LABEL:table:lmms. As we can see, ShareGPT4V achieves the best performance in terms of PerceptionScore, followed by LLaMA-Adapter V2, which is used as the Image-to-Text models in [74]. Additionally, we note that there is study [86] argue that InstructBLIP [14] exhibits state-of-the-art (SOTA) performance in image-to-text generation. However, in our experiment, we found that InstructBLIP is unsuitable for handling geospatial coordinates as prompt content. Incorporating such coordinates significantly degrades the quality of the generated text, leading to poor content output. Due to these limitations, we have excluded it from our comparative analysis.
Table 3: Perception score of representative LMMs.
@@
LLaVA-v1.5 [48] 0.674
mPLUG-Owl [77] 0.662
LLaMA-Adapter V2 [20] 0.698
MiniGPTv2 [10] 0.693
ShareGPT4V [11] 0.714
Appendix H More Experimental Analysis
H.1 More Analysis for Text Calibration
In our paper, we selected the state-of-the-art LMM ShareGPT4V, which demonstrates performance comparable to GPT4V in image captioning. In order to ascertain the efficacy of ShareGPT4V, we computed the distribution of PerceptionScore for descriptions generated by ShareGPT4V, as depicted in Figure 14. Notably, only less than 0.02% of generated texts scored below 0.5, thereby evidencing the high quality of the descriptions produced.
Table 4: UrbanVLP performance utilizing mPLUG-Owl with and without Text Calibration. mPO is an abbreviation for mPLUG-Owl, and TC stands for Text Calibration.
Model Carbon Population GDP
$\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE
UrbanVLP_mPO_wo TC 0.687 0.551 0.544 0.643 0.580 0.458 0.494 0.655 0.447
UrbanVLP_mPO 0.714 0.538 0.473 0.683 0.557 0.430 0.506 0.640 0.435
UrbanVLP 0.769 0.477 0.369 0.714 0.523 0.411 0.537 0.684 0.416
As a result, the experimental results comparing with and without calibration do not adequately reflect the effectiveness of our calibration method. To further demonstrate the effectiveness of our method, we select mPLUG-Owl [77], one of the representative LMMs compared in our paper, with the worst performance. After calibration, 15.3% of the text with perceptionscores below 0.5 was filtered out. The results validate the effectiveness of our calibration method, and indicate that more relevant and noise-free textual information could align better with image features, leading to a more coherent and meaningful visual representation.
H.2 Analysis about Time Complexity
Modeling and research in urban areas can provide valuable references and assistance to policymakers, builders, and citizens. Therefore, practical implementation should be considered. We conduct time complexity analysis using FPS (Frames Per Second), a common indicator for inference speed in computer vision area. We report FPS for satellite image in Figure 16(b). In fact, introducing street view features for better and finer-grained effects will inevitably sacrifice a portion of the time efficiency. However, compared to [41]’s graph structure inference, our linear probing setting ensures the minimum time overhead, also with better performance. For making model decisions during deployment, to pursue the trade-off between performance and speed, UrbanVLP is the optimal choice.
Figure 16: (a) Hyperparameter for loss combination. (b) FPS comparison of different mdoels. (c) Expetiments for maximum of the number of street-view images.
H.3 Analysis about Hyperparameters
The optimal hyperparameters for Equation 7 are actually determined by grid search. In Figure 16(a), we present some results of different combinations of hyperparameter values . It can be seen that the settings we chose in the paper ($\alpha$=0.5, $\beta$=0.5) yield the best experimental results.
We also study the maximum value of street-view images during training, which is set as 25 in our paper. As is shown in Figure 16(c), as the number of street-view images increases, the R² value continuously rises, reaching its peak at 25. Therefore, our setting is the optimal choice.
Appendix I Practicality
Figure 17: UrbanVLP system.